Suppose that a monopoly faces inverse market demand function as

P = 70−2Q,

and its marginal cost function is

MC = 40 – Q.

Please answer the following two questions:

a. What should be the monopoly’s profit-maximizing output?

b. What is the monopoly’s price?
Answer #1

P=70-2Q

MC=40-Q

MR=70-4Q

Setting MR=MC

70-4Q=40-Q

70-40=-Q+4Q

30=3Q

a) Q=30/3=10

b) P=70-2(10) = 50